# pFedMMA: Personalized Federated Fine-Tuning with Multi-Modal Adapter for Vision-Language Models

### v2

@@ 4.3 Ablation Study
Dimension of the Shared Layer. Table 5 (bottom-left) reports the average accuracies over four datasets and five shot settings. As shown, using a larger 128-dimensional representation yields slightly better performance than 32 dimensions. However, to keep the number of trainable parameters low, we consistently adopt the 32-dimensional setting throughout the paper. Detailed results are provided in Tables 11 and 12 in the Appendix.
@@
Adapting Variant Options for PFL. We evaluate the effectiveness of different design choices of MMA in personalized federated learning. In Table 5 (bottom-right), we compare two alternative baselines: treating all adapters as global (Baseline 1) and using the shared adapter as the personalized component while treating the up- and down-projection adapters as global (Baseline 2). As shown, pFedMMA achieves significantly higher local accuracy than both baselines. Moreover, it achieves superior base and novel performance compared to state-of-the-art prompt learning methods, as shown earlier, underscoring its ability to strike a strong balance between personalization and generalization.
@@ 4.4 Learning Curves
To further analyze the convergence behavior of pFedMMA, we plot the average local accuracy over communication rounds across five different shot settings in Fig. 4. As shown, pFedMMA consistently attains high accuracy and converges faster than the baselines. These results highlight its ability to effectively balance personalization and generalization, ensuring both rapid and reliable convergence. Detailed results are provided in Fig. 5 in the Appendix.
@@ Appendix A Pipelines of the Proposed Algorithm
12:  At server: ${\bm{W}}_{js}^{t+1}=\frac{1}{N}\sum_{i=1}^{N}{\bm{W}}_{js,i}^{t,E},\quad j\in\{\ell,\cdots,L\}$
@@ B.1 Personalized Federated Learning
Personalized Federated Learning (PFL) has emerged as a pivotal research direction to address the limitations of conventional federated learning McMahan et al. (2017) when faced with heterogeneous client data. Unlike standard FL, which learns a single global model, PFL aims to produce tailored models for individual clients, thus better coping with statistical and systemic heterogeneity Tan et al. (2022); Kulkarni et al. (2020). Several personalization strategies have been proposed, including local fine-tuning Mansour et al. (2020); Tan et al. (2022); Wang et al. (2019), regularization-based optimization Li et al. (2020; 2021b); T Dinh et al. (2020), and parameter decomposition into shared and client-specific components Arivazhagan et al. (2019); Oh et al. ; Collins et al. (2021). Other methods pursue clustering of clients to exploit latent similarities Huang et al. (2021); Zhang et al. ; Sattler et al. (2020); Ziad et al. (2024), or leverage attention mechanisms and adaptive layers Liang et al. (2020); Li et al. (2023); Sun et al. (2023). To further improve adaptability, techniques like FedBN Li et al. and PartialFed Sun et al. (2021) address feature shift via local normalization or selective personalization. Hybrid global-local learning approaches have also been developed Deng et al. (2020); Chen & Chao (2022). FedOT Farnia et al. (2022) proposes learning optimal transport maps that align local distributions to a shared probability space, enabling a global classifier to be trained more effectively; personalization is achieved by composing this shared model with each client’s transport map. While these approaches have demonstrated success, they typically center on traditional ML architectures and do not yet fully leverage the potential of large pre-trained models, such as vision-language or foundation models, for personalization.
@@ B.2 Federated Prompt Learning for VLMs
Federated Prompt Learning (FPL) extends the flexibility of prompt tuning for adapting large pre-trained models such as CLIP Radford et al. (2021) to the FL settings, enabling efficient and personalized downstream task adaptation across decentralized clients. Early works like CoOp Zhou et al. (2022b) and CoCoOp Zhou et al. (2022a) laid the foundation by introducing learnable continuous prompt vectors, which sparked interest in federated extensions. PromptFL Guo et al. (2023b) and FedPrompt Zhao et al. (2022) introduced FL-style prompt aggregation, performing FedAvg McMahan et al. (2017) over client-specific prompt updates. FedPR Feng et al. (2023) explores visual prompt learning within the null space of global prompts for MRI reconstruction, while FedAPT Su et al. (2022) focuses on domain-adaptive prompt tuning for cross-domain image classification. To enhance personalization, pFedPrompt Guo et al. (2023a) introduces a non-parametric attention module over local few-shot memory, and pFedPG Yang et al. (2023) and FedTPG Qiu et al. design server-side prompt generators to issue personalized prompts to each client. FedCLIP Lu et al. integrates attention-based adapters to better exploit the pre-trained model’s knowledge. Furthermore, FedOTP Li et al. (2024) leverages Optimal Transport to align global and local prompts, and FedPGP Cui et al. (2024) utilizes prompt-wise contrastive losses to better capture diverse category-level traits across clients. Recently, pFedMoAP Luo et al. (2025) rethinks prompt sharing by treating pre-aggregated prompts from other clients as non-local experts in a Mixture-of-Experts framework, enabling effective personalization via a lightweight, attention-based gating mechanism. Theoretical analysis of FPL Pan et al. (2024) provides deeper understanding of its convergence properties.
@@ C.1 Dataset Setup
Flowers102 Nilsback & Zisserman (2008) Fine-grained flowers recognition 102 10 100% 4,093 2,463 1 Pathological
@@ E.1 Comparison
Table 17: Comparison of computation, communication, and accuracy for five personalized federated learning methods under CLIP ViT-B/16. Columns report the number of local trainable parameters, per-round communicated parameters (upload/down), end-to-end training time, peak GPU memory, average local accuracy, and average harmonic-mean (HM) accuracy; notation follows Table 16.

### v3

@@ 4.3 Ablation Study & Hyperparameter Sensitivity
Dimension of the Shared Layer. Table 5 (bottom-left) reports the average accuracies over four datasets and five shot settings. As shown, using a larger 128-dimensional representation yields slightly better performance than 32 dimensions. However, to keep the number of trainable parameters low, we consistently adopt the 32-dimensional setting throughout the paper. Detailed results are provided in Tables 13 and 14 in the Appendix.
@@
Adapting Variant Options for Personalization. We evaluate the effectiveness of different design choices of MMA in personalized federated learning. In Table 5 (bottom-right), we compare two alternative baselines: treating all adapters as global (Baseline 1) and using the shared adapter as the personalized component while treating the up- and down-projection adapters as global (Baseline 2). As shown, pFedMMA achieves significantly higher local accuracy than both baselines. Moreover, it achieves superior base and novel performance compared to state-of-the-art prompt learning methods, as shown earlier, underscoring its ability to strike a strong balance between personalization and generalization.
Table 6: Comparison of FL aggregation variants (vision-only, text-only, and both-sides) for the shared adapter.
Methods Local Base Novel HM
Vision Only $95.81$ $71.19$ $76.07$ $79.31$
Text Only $95.99$ $71.19$ $76.13$ $79.38$
Both Vision & Text $95.99$ $71.24$ $76.10$ $79.39$
pFedMMA (Ours) 96.14 71.78 76.17 79.70
Adapting Variant Options for FL Aggregation. We next ablate how the shared adapter is aggregated across clients to localize the main information-sharing channel. In Table 6, we compare three variants that differ in which modality-specific shared block is federated: Vision Only, where only the vision-side shared block is aggregated; Text Only, where only the text-side shared block is aggregated; and Both Vision & Text, where separate shared blocks for each modality are aggregated simultaneously. These variants achieve very similar local accuracy, with small but consistent differences in base, novel, and HM: aggregating text-only or both modalities yields a slight edge in HM over aggregating vision-only. Our full pFedMMA, which uses a single multi-modal shared projection rather than two separate modality-specific ones, further improves local, base, novel, and HM over all three variants, suggesting that tying the modalities through a unified shared adapter provides a slightly stronger and more stable information-sharing mechanism without harming personalization.
@@ 4.4 Learning Curves
To further analyze the convergence behavior of pFedMMA, we plot the average local accuracy over communication rounds across five different shot settings in Fig. 4. As shown, pFedMMA consistently attains high accuracy and converges faster than the baselines. Detailed results are provided in Fig. 5 in the Appendix.
@@
4.5 Training Cost Analysis
Table 7 summarizes the computational and communication costs together with accuracy. PromptFL has the smallest footprint (8,192 trainable; 8,192/8,192 per round) but—most importantly—shows a marked drop in local accuracy (88.93%), indicating weaker personalization under heterogeneity. FedPGP increases local trainables to 12,416 without extra communication but incurs the highest memory (13,374 MiB) and slower training, and its HM accuracy (79.09%) lags its strong local accuracy (95.38%). FedOTP doubles prompt capacity (16,384 trainables; same 8,192/8,192 communication) and attains very high local accuracy (97.34%) but suffers extremely low HM accuracy (31.08%), suggesting poor cross-client generalization. pFedMoAP adds a local gating module, raising local trainables (74,240) and the per-round download (73,728) while achieving the shortest training time (902 s) and strong local accuracy (97.89%). pFedMMA (ours) communicates only the shared adapter blocks (3,072/3,072) while keeping 248,832 parameters local, yielding the best HM accuracy (84.15%) and competitive local accuracy (97.17%), thus offering the most favorable accuracy–communication trade-off. Please refer to Appendix E for additional details.
Table 7: Comparison of computation, communication, and accuracy for five personalized federated learning methods under CLIP ViT-B/16.
PromptFL Guo et al. (2023a) 8,192 8,192 / 8,192 1,645 5,116 88.93 83.09
FedPGP Cui et al. (2024) 12,416 8,192 / 8,192 3,980 13,374 95.38 79.09
FedOTP Li et al. (2024) 16,384 8,192 / 8,192 1,328 3,014 97.34 31.08
pFedMoAP Luo et al. (2025) 74,240 8,192 / 73,728 902 3,108 97.89 71.05
pFedMMA (Ours) 248,832 3,072 / 3,072 2,175 4,634 97.17 84.15
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Science Foundation under Grant 2419982, Grant 2342253, and Grant 2236483.
@@ Appendix A Pipelines of the Proposed Algorithm
12:  At server: ${\bm{W}}_{js}^{t+1}=\sum_{i=1}^{N}p_{i}{\bm{W}}_{js,i}^{t,E},\quad j\in\{\ell,\cdots,L\}$
@@ B.1 Personalized Federated Learning
Personalized Federated Learning (PFL) has emerged as a pivotal research direction to address the limitations of conventional federated learning McMahan et al. (2017) when faced with heterogeneous client data. Unlike standard FL, which learns a single global model, PFL aims to produce tailored models for individual clients, thus better coping with statistical and systemic heterogeneity Tan et al. (2022); Kulkarni et al. (2020). Several personalization strategies have been proposed, including local fine-tuning Mansour et al. (2020); Tan et al. (2022); Wang et al. (2019), regularization-based optimization Li et al. (2020; 2021b); T Dinh et al. (2020), and parameter decomposition into shared and client-specific components Arivazhagan et al. (2019); Oh et al. ; Collins et al. (2021). Other methods pursue clustering of clients to exploit latent similarities Huang et al. (2021); Zhang et al. ; Sattler et al. (2020); Ziad et al. (2024), or leverage attention mechanisms and adaptive layers Liang et al. (2020); Li et al. (2023a); Sun et al. (2023). To further improve adaptability, techniques like FedBN Li et al. and PartialFed Sun et al. (2021) address feature shift via local normalization or selective personalization. Hybrid global-local learning approaches have also been developed Deng et al. (2020); Chen and Chao (2022). FedOT Farnia et al. (2022) proposes learning optimal transport maps that align local distributions to a shared probability space, enabling a global classifier to be trained more effectively; personalization is achieved by composing this shared model with each client’s transport map. While these approaches have demonstrated success, they typically center on traditional ML architectures and do not yet fully leverage the potential of large pre-trained models, such as vision-language or foundation models, for personalization.
@@ B.2 Federated Prompt Learning for VLMs
Federated Prompt Learning (FPL) extends the flexibility of prompt tuning for adapting large pre-trained models such as CLIP Radford et al. (2021) to the FL settings, enabling efficient and personalized downstream task adaptation across decentralized clients. Early works like CoOp Zhou et al. (2022b) and CoCoOp Zhou et al. (2022a) laid the foundation by introducing learnable continuous prompt vectors, which sparked interest in federated extensions. PromptFL Guo et al. (2023a) and FedPrompt Zhao et al. (2022) introduced FL-style prompt aggregation, performing FedAvg McMahan et al. (2017) over client-specific prompt updates. FedPR Feng et al. (2023) explores visual prompt learning within the null space of global prompts for MRI reconstruction, while FedAPT Su et al. (2022) focuses on domain-adaptive prompt tuning for cross-domain image classification. To enhance personalization, pFedPrompt Guo et al. (2023b) introduces a non-parametric attention module over local few-shot memory, and pFedPG Yang et al. (2023) and FedTPG Qiu et al. design server-side prompt generators to issue personalized prompts to each client. FedCLIP Lu et al. integrates attention-based adapters to better exploit the pre-trained model’s knowledge. Furthermore, FedOTP Li et al. (2024) leverages Optimal Transport to align global and local prompts, and FedPGP cui2024harmonizing utilizes prompt-wise contrastive losses to better capture diverse category-level traits across clients. Recently, pFedMoAP Luo et al. (2025) rethinks prompt sharing by treating pre-aggregated prompts from other clients as non-local experts in a Mixture-of-Experts framework, enabling effective personalization via a lightweight, attention-based gating mechanism. Theoretical analysis of FPL Pan et al. (2024) provides deeper understanding of its convergence properties.
@@ B.3 Efficient Transfer Learning for VLMs
B.4 Federated Out-of-Distribution and Domain Generalization
A complementary line of work studies federated domain generalization (FedDG), where the goal is to train a single global model that generalizes to unseen target domains under client heterogeneity. FedDAT Chen et al. (2024) tackles multi-modal heterogeneous FL for foundation vision-language models via a dual-adapter teacher and mutual knowledge distillation, improving global performance across diverse vision-language tasks under domain shift. PLAN Gong et al. (2024) introduces a FedDG framework for pre-trained vision-language models based on visual and textual prompt learning and attention-based prompt aggregation, explicitly using a leave-one-domain-out protocol to adapt a global CLIP-style model to unseen domains. Other recent methods similarly design adapter- or prompt-based FedDG algorithms to enhance out-of-domain robustness of federated foundation models Li et al. (2023b); Yang et al. (2025); Lu et al. (2023). In contrast, our work is formulated in the personalized federated learning (PFL) setting, where each client maintains its own model (or adapter) and we explicitly optimize the personalization–generalization trade-off; consequently, we treat FedDG methods as complementary rather than direct baselines, and instead compare against personalized prompt- and adapter-based methods, while showing that our approach achieves comparable personalized performance and substantially stronger cross-domain generalization within this PFL protocol.
@@ C.1 Dataset Setup
Flowers102 Nilsback and Zisserman (2008) Fine-grained flowers recognition 102 10 100% 4,093 2,463 1 Pathological
@@ D.4 Ablation Study
Table 18: Comparison of FL aggregation variants (vision-only, text-only, and both-sides) for the shared adapter.
Shots Layers Average on 4 datasets SUN397 Flowers102 DTD Food101
16 Vision Only $95.81$ $71.19$ $76.07$ $79.31$ $94.02$ $70.73$ $76.08$ $79.12$ $95.79$ $69.62$ $76.31$ $79.14$ $97.31$ $55.22$ $61.10$ $67.04$ $96.13$ $89.17$ $90.80$ $91.94$
Text Only $95.99$ $71.19$ $76.13$ $79.38$ $94.20$ $70.80$ $76.10$ $79.20$ $96.40$ $69.53$ $76.30$ $79.24$ $97.08$ $55.28$ $61.32$ $67.12$ $96.27$ $89.17$ $90.80$ $91.98$
Both Vision & Text $95.99$ $71.24$ $76.10$ $79.39$ $94.23$ $70.85$ $76.11$ $79.23$ $96.03$ $69.61$ $76.25$ $79.17$ $97.27$ $55.32$ $61.24$ $67.13$ $96.44$ $89.17$ $90.80$ $92.03$
pFedMMA (Ours) $96.14$ $71.78$ $76.17$ $79.70$ $94.06$ $70.99$ $76.37$ $79.34$ $95.58$ $71.54$ $76.00$ $79.79$ $97.45$ $55.44$ $61.55$ $67.35$ $97.45$ $89.15$ $90.77$ $92.32$
Table 19: Top-1 accuracy (%) of different methods across 7 datasets in the 16-shot setting using AdamW optimizer.
Average on 7 datasets SUN397 Flowers102 DTD
PromptFL Guo et al. (2023a) $86.80$ $86.87$ $79.36$ $84.19$ $77.44$ $77.42$ $73.63$ $76.12$ $89.69$ $89.74$ $73.62$ $83.62$ $76.16$ $75.81$ $54.11$ $66.96$
FedPGP Cui et al. (2024) $97.10$ $63.53$ $67.19$ $73.31$ $95.57$ $41.24$ $49.61$ $54.68$ $99.57$ $47.37$ $57.42$ $61.77$ $95.46$ $48.77$ $44.69$ $56.23$
pFedMoAP Luo et al. (2025) $97.96$ $61.35$ $67.59$ $72.63$ $96.17$ $33.01$ $36.36$ $43.99$ $99.86$ $36.36$ $51.83$ $52.81$ $96.48$ $50.90$ $46.06$ $58.00$
pFedMMA (Ours) $96.29$ $77.90$ $81.65$ $84.58$ $93.78$ $71.17$ $76.51$ $79.40$ $95.79$ $70.93$ $76.85$ $79.89$ $92.82$ $56.91$ $61.50$ $67.26$
OxfordPets Caltech101 Food101 UCF101
PromptFL Guo et al. (2023a) $96.19$ $96.01$ $96.64$ $96.28$ $97.25$ $98.13$ $92.90$ $96.04$ $90.73$ $90.76$ $91.15$ $90.88$ $80.12$ $80.20$ $73.50$ $77.81$
FedPGP Cui et al. (2024) $98.78$ $86.27$ $94.47$ $92.88$ $99.83$ $87.82$ $88.09$ $91.59$ $95.96$ $78.62$ $78.67$ $83.68$ $94.56$ $54.64$ $57.36$ $64.78$
pFedMoAP Luo et al. (2025) $99.90$ $78.13$ $91.76$ $89.00$ $99.94$ $94.62$ $92.47$ $95.58$ $97.60$ $71.43$ $84.37$ $83.11$ $95.76$ $64.98$ $70.26$ $74.88$
pFedMMA (Ours) $99.08$ $89.84$ $96.74$ $95.05$ $100.0$ $97.10$ $94.18$ $97.04$ $97.24$ $89.42$ $90.75$ $92.35$ $95.31$ $69.90$ $75.02$ $78.68$
@@ E.1 Comparison
Table 21: Comparison of computation, communication, and accuracy for five personalized federated learning methods under CLIP ViT-B/16. Columns report the number of local trainable parameters, per-round communicated parameters (upload/down), end-to-end training time, peak GPU memory, average local accuracy, and average harmonic-mean (HM) accuracy.